OpenStudy (anonymous):

What is the approximate mass of 6.022x10^23 12C atoms?

OpenStudy (anonymous):



OpenStudy (anonymous):

@somy could you help me with this?

OpenStudy (anonymous):

What does the 12 C means?

OpenStudy (somy):

12 is mass number

OpenStudy (somy):

we write 12 C because there are isotopes of Carbon also so by writing 12 we show which carbon isotope it is

OpenStudy (somy):

don't mind that they are asking for mass of an atom

OpenStudy (somy):

after finding mass of 1 atom u can find mass of 1 mole of atoms

OpenStudy (somy):

atom consists of protons, neutrons and electrons

OpenStudy (anonymous):

I know to find the mass of C is (1atom of C * 12g of C)/6.022x10^23 atoms of C

OpenStudy (anonymous):

That would be the weight of 1 atom of Carbon

OpenStudy (somy):

mole= mass/ Mr mass= mole*Mr mole is 1 because 1 mole has 6.022x10^23 atoms Mr is 12

OpenStudy (anonymous):

What is Mr?

OpenStudy (somy):

molecular mass

OpenStudy (anonymous):

How do I solve this then? I keep trying and I don't get the right answer.

OpenStudy (somy):

what is the right answer?

OpenStudy (somy):

7.2264x10^24

OpenStudy (anonymous):

yes, that's right. How do I obtain that?

OpenStudy (somy):

1 atom --- 12 g/mol 6.022x10^23 atoms --- X X= (6.022x10^23) * 12

OpenStudy (anonymous):

x= 1.9927x10^-23 What do I do after I find this?

OpenStudy (anonymous):

oh wait...

OpenStudy (anonymous):

What's the point of finding the mass of 1 atom?

OpenStudy (somy):

lol u are making mistake

OpenStudy (somy):

write 6.022x10^23 in bracket in ur calculator like this (6.022x10^23)

OpenStudy (somy):

then multiply by 12

OpenStudy (anonymous):

yes I know that part. I am wondering about the weight of 1 atom

OpenStudy (somy):

well look mass of 1 atom would be its molecular mass

OpenStudy (anonymous):

I am asked to find the mass of 6.022x10^23 C atoms.

OpenStudy (somy):

but they are indirectly asking u for mass of 1 mole of 12C

OpenStudy (anonymous):

1atom C*1mole/6.022x10^23 atom * 12g/1mol

OpenStudy (somy):

1 atom of C has a mass of 12 amu 1 mole of C has 6.022x10^23 atoms so if 1 atom has mass of 12 amu how many amu would be mass of 6.022x10^23 atoms?

OpenStudy (anonymous):

6.022x10^23 * 12 o.o

OpenStudy (anonymous):

the amu unit confuses meee D;

OpenStudy (anonymous):

since is in amu and not in mol, it is not 1mol C * 12g C, instead is 6.022x10^23 C amu * 12amu

OpenStudy (somy):

yes

OpenStudy (somy):

my god u are confusing me 6.022x10^23 atoms ---- u need to find amu (amu is molecular mass unit)

OpenStudy (anonymous):

but but you said yes =.=''

OpenStudy (somy):

u have mass of 1 atom and its 12 now they gave you 6.022x10^23 atoms so if u know mass of 1 atom u can find mass of 6.022x10^23 atoms right?

OpenStudy (somy):

alright look u have 1 apple and mass of is 30 grams and then u have same apple but now 50 of them

OpenStudy (somy):

if 1 apple is 30 grams then 50 apples would be 30*50 = 1500 grams

OpenStudy (anonymous):

How would you do it in the choo choo train?

OpenStudy (somy):

lol jokes aside do u understand? after doing this (6.022x10^23) *12 you'll get 7.2264x10^24

OpenStudy (anonymous):

it wasn't a joke =.=. that's how they call it here lol.

OpenStudy (anonymous):

yes I understand that, I am trying to figure it out starting from the 6.022x10^23 atoms. with the train.

OpenStudy (anonymous):

What do I multiply it by or divide to cancel atoms

OpenStudy (somy):

u want to get the unit?

OpenStudy (anonymous):

Yes, step by step. I want to understand the mathematic way of getting the answer.

OpenStudy (somy):

\[mole= \frac{ mass }{molecular mass }\]

OpenStudy (somy):

\[mass= mole * molecular mass\]

OpenStudy (somy):

now units unit of mole is mole unit of molecular mass is g/mole \[mole *\frac{ gram }{ mole }\]

OpenStudy (anonymous):

since one side is in molecular mass (amu) then the mole in order to have the right unit must be in atoms.

OpenStudy (somy):

moles cancel out and you are left with grams